Skip to main contentSkip to solution

In a binomial distribution, the sum of its mean and variance is 1.8. If the event was conducted 5 times, then the probability of 3 successes is

Solution

✅ Correct Option: 1

With n=5n=5, mean ++ variance =np+npq=5p(2−p)=1.8=np+npq=5p(2-p)=1.8, so 5p2−10p+1.8=05p^2-10p+1.8=0 and p=0.2p=0.2. Then P(3)=(53)(0.2)3(0.8)2=10×0.008×0.64=0.0512P(3)=\binom53(0.2)^3(0.8)^2=10\times0.008\times0.64=0.0512.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question